Two pans of a balance are 22.4 cm apart.
The fulcrum of the balance has been shifted
1.01 cm away from the center by a dishonest
shopkeeper.
By what percentage is the true weight of the
goods being marked up by the shopkeeper?
Assume the balance has negligible mass.
Answer in units of %. Don't round answer.
1
Expert's answer
2020-01-20T05:21:20-0500
"Wd=W'd'"

"W(22.4-1.01)=W'(22.4+1.01)"

"W'=0.9137W"

Now we can find the percentage change as


"\frac{W-W'}{W}100\%=\frac{W-0.9137W}{W}100\%=8.63\%"